OpenStudy (anonymous):

solve for y: 4x-2y=12

OpenStudy (anonymous):

4x - 2y = 12 4(0) - 2y = 12 2y = 12 12 / 2 = 6 y = 6 do you understand ?

OpenStudy (anonymous):

thank you and yes

OpenStudy (anonymous):

your welcome, and thanks for the medal as well :)